# Robust Watermarking Using Generative Priors Against Image Editing: From Benchmarking to Advances

### v1

@@ Abstract
Current image watermarking methods are vulnerable to advanced image editing techniques enabled by large-scale text-to-image models. These models can distort embedded watermarks during editing, posing significant challenges to copyright protection. In this work, we introduce W-Bench, the first comprehensive benchmark designed to evaluate the robustness of watermarking methods against a wide range of image editing techniques, including image regeneration, global editing, local editing, and image-to-video generation. Through extensive evaluations of eleven representative watermarking methods against prevalent editing techniques, we demonstrate that most methods fail to detect watermarks after such edits. To address this limitation, we propose VINE, a watermarking method that significantly enhances robustness against various image editing techniques while maintaining high image quality. Our approach involves two key innovations: (1) we analyze the frequency characteristics of image editing and identify that blurring distortions exhibit similar frequency properties, which allows us to use them as surrogate attacks during training to bolster watermark robustness; (2) we leverage a large-scale pretrained diffusion model SDXL-Turbo, adapting it for the watermarking task to achieve more imperceptible and robust watermark embedding. Experimental results show that our method achieves outstanding watermarking performance under various image editing techniques, outperforming existing methods in both image quality and robustness. Code is available at https://github.com/Shilin-LU/VINE
@@ 2 Related Work
Robust watermarking. Image watermarking has long been studied for purposes such as tracking and protecting intellectual property (Al-Haj, 2007; Cox et al., 2007; Navas et al., 2008). Recently, deep learning-based methods (Bui et al., 2023; Chen & Li, 2024; Fang et al., 2022; 2023; Jia et al., 2021; Kishore et al., 2021; Luo et al., 2020; 2024; Ma et al., 2022; Tancik et al., 2020; Wu et al., 2023; Zhu, 2018; Zhang et al., 2019; 2021) have demonstrated competitive robustness against a wide range of transformations. However, these methods remain vulnerable to image editing powered by large-scale generative models. Three recent studies—EditGuard (Zhang et al., 2024d), Robust-Wide (Hu et al., 2024), and JigMark (Pan et al., 2024)—have begun to develop watermarking models that are robust to such image editing. However, although EditGuard, which replaces the unedited regions of an edited image with corresponding areas of the unedited watermarked version, demonstrates good performance, its application in real-world scenarios could be ineffective if the detector has no access to the unedited watermarked image. In our benchmark, we assume that the detector has access only to the edited image without any additional information. JigMark employs contrastive learning to train a watermark encoder and a classifier that determines if an image is watermarked, but it does not support decoding free-form messages. Robust-Wide incorporates Instruct-Pix2Pix into the noise layer using gradient truncation, but its generalization to other editing models could be less effective. Moreover, neither JigMark nor Robust-Wide has been open-sourced.
@@ 3.1 Frequency nature of image editing
To develop a robust watermarking model against image editing, it should learn to embed information into the low-frequency bands. To identify effective surrogate attacks, we explore various image distortions, denoted as $\mathcal{T}(\cdot)$, that resemble image editing to some extent. Both image editing and distortion methods preserve the overall image layout and most of the content, although image distortions typically result in lower perceptual quality. Notably, as shown in Figure 3, certain blurring distortions (e.g., pixelation and defocus blur) exhibit a similar trend to image editing, where patterns in the low-frequency bands are less affected than those in the mid- and high-frequency bands. In contrast, widely used distortions, such as JPEG compression and saturation, do not display this behavior. Since these blurring distortions are computationally efficient, we incorporate them with varying severity levels into our noise layers during training. This encourages the model to embed information within the low-frequency bands (see the frequency patterns of each watermarking method in Appendix B). Consequently, the robustness against image editing is enhanced, as demonstrated by the ablation study in Table 2. The complete set of distortions applied in the noise layer includes common distortions such as saturation, contrast, brightness adjustments, JPEG compression, Gaussian noise, shot noise, impulse noise, and speckle noise to counteract image degradation caused by transmission. Additionally, we incorporate various blurring distortions—such as pixelation, defocus blur, zoom blur, Gaussian blur, and motion blur—to resist image editing.
@@ 3.3 Objective function and training strategy
Training strategy. In the first training stage, we prioritize the watermark extraction loss by setting $\alpha$ to 10 and $\beta_{\text{MSE}}$, $\beta_{\text{LPIPS}}$, $\beta_{\text{GAN}}$ each to 0.01. To preserve the generative prior, the UNet and VAE decoder of the SDXL-Turbo, along with the added zero-convolution layers, are frozen. Once the bit accuracy exceeds 0.85, we transition to the second stage, unfreezing all parameters for further training. At this point, the loss weighting factors are adjusted to $\alpha=1.5$, $\beta_{\text{MSE}}=2.0$, $\beta_{\text{LPIPS}}=1.5$, and $\beta_{\text{GAN}}=0.5$. The model after the first two stages serves as our base model, dubbed VINE-B. We then fine-tune VINE-B in the third stage by incorporating Instruct-Pix2Pix (Brooks et al., 2023), a representative instruction-driven image editing model, into our noise layer. Gradients are backpropagated through a straight-through estimator (Bengio et al., 2013). Note that it cannot converge if it is applied directly during the early training stages. The fine-tuned model is referred to as VINE-R. Additional implementation details are provided in Appendix E.
Resolution scaling. Different watermarking models are typically trained using a fixed input resolution, limiting them to accepting only fixed-resolution inputs during testing. However, in practical applications, supporting watermarking at the original resolution is crucial to preserve the input image quality. Bui et al. (2023) propose a method (detailed in Appendix C.1) to adapt any watermarking model to handle arbitrary resolutions without compromising the quality of the watermarked images and their inherent robustness, as demonstrated in Appendices C.2 and C.3. In our experiments, we apply this resolution scaling method to all methods, enabling them to operate at a uniform resolution of $512\times 512$, which is compatible with image editing models.
@@ 4.1 Image editing methods
Image regeneration. Image regeneration involves perturbing an image into a noisy version and then reconstructing it. The perturbing process can be either stochastic or deterministic. In the stochastic method (Meng et al., 2021; Nie et al., 2022; Zhao et al., 2023b), random Gaussian noise is introduced to the image, with the noise level typically controlled by a timestep $t_{s}$ using common noise schedulers such as VP (Ho et al., 2020), VE (Song & Ermon, 2019; Song et al., 2020b), FM (Lipman et al., 2022; Liu et al., 2022a), and EDM (Karras et al., 2022). The diffusion model then denoises the noisy image starting from timestep $t_{s}$ to produce a clean image. In contrast, the deterministic method (Mokady et al., 2022; Song et al., 2020a; Wallace et al., 2022), also known as image inversion, utilizes a diffusion model to deterministicly invert a clean image into a noisy version through multiple sampling steps $n_{d}$. The image is then reconstructed by applying the same sampling methods over the same number of sampling steps $n_{d}$. We employ the widely used VP scheduler for stochastic regeneration, testing noise timesteps $t_{s}$ from 60 to 240 in increments of 20. For deterministic regeneration, we utilize the fast sampler DPM-solver (Lu et al., 2022a; b) and evaluate sampling steps $n_{d}$ of 15, 25, 35, and 45. See Appendix F.1 for the effects of image regeneration on images.
@@ 4.2 Experimental setup
Table 1: Comparison of watermarking performance in terms of watermarked image quality and detection accuracy across various image editing methods. Quality metrics are averaged over 10,000 images, and the TPR@0.1%FPR for each specific editing method is averaged over 5,000 images. The best value in each column is highlighted in bold, and the second best value is underlined. Abbreviations: Cap = Encoding Capacity; Sto = Stochastic Regeneration; Det = Deterministic Regeneration; Pix2Pix = Instruct-Pix2Pix; Ultra = UltraEdit; Magic = MagicBrush; CtrlN = ControlNet-Inpainting; SVD = Stable Video Diffusion.
@@
Datasets. We train VINE using the OpenImage dataset (Kuznetsova et al., 2020) at a resolution of $256\times 256$. The training details are provided in Appendix E. For evaluation, we randomly sample 10,000 instances from the UltraEdit dataset (Zhao et al., 2024c), each containing a source image, an editing prompt, and a region mask. The images in UltraEdit dataset are photographs sourced from datasets such as COCO (Lin et al., 2014), Flickr (Young et al., 2014), and ShareGPT4V (Chen et al., 2023). Of these 10,000 samples, 1,000 are allocated for stochastic regeneration, another 1,000 for deterministic regeneration, and 1,000 for global editing. 5,000 samples are designated for local editing. These consist of five sets, each containing 1,000 images, which are edited using mask sizes covering 10–20%, 20–30%, 30–40%, 40–50%, and 50–60% of the total image area. Additionally, we include 1,000 samples for image-to-video generation and 1,000 for testing conventional distortion, thereby completing the total evaluation set.
Baselines. We compare VINE with eleven watermark baselines, all utilizing their officially released checkpoints. These baselines include MBRS (Jia et al., 2021), CIN (Ma et al., 2022), PIMoG (Fang et al., 2022), RivaGAN (Zhang et al., 2019), SepMark (Wu et al., 2023), TrustMark (Bui et al., 2023), DWTDCT (Al-Haj, 2007), DWTDCTSVD (Navas et al., 2008), SSL (Fernandez et al., 2022), StegaStamp (Tancik et al., 2020), and EditGuard (Zhang et al., 2024d). Although the baselines were trained at different fixed resolutions (as detailed in Appendix C), we apply resolution scaling (described in Section 3.3) to standardize all of them to a uniform resolution of $512\times 512$. This standardization does not compromise their robustness, as demonstrated in Appendix C.3.
Metrics. We evaluate the imperceptibility of watermarking models using standard metrics, including PSNR, SSIM, LPIPS (Zhang et al., 2018), and FID (Parmar et al., 2022). For watermark extraction, it is essential to strictly control the false positive rate (FPR), as incorrectly labeling non-watermarked images as watermarked can be detrimental—a concern often overlooked in previous studies. Neither high bit accuracy nor AUROC alone guarantees a high true positive rate (TPR) at low FPR levels. Thus, we primarily focus on TPR@0.1%FPR and TPR@1%FPR as our main metrics. Accordingly, both the watermarked and original images are fed into watermark decoders for evaluation. Additionally, we also provide bit accuracy and AUROC for reference. Note that all reported baseline bit accuracies do not include error correction methods, such as BCH (Bose & Ray-Chaudhuri, 1960), which can be applied to all watermarking models.
@@ 4.3 Benchmarking results and analysis
Table 1 summarizes the overall evaluation results. As discussed in Section 4.2, we report TPR@0.1%FPR as the primary metric, with additional metrics provided in Figure 5. Each reported TPR@0.1%FPR value is averaged across a total of $m\times$1,000 images, where $m$ represents the number of difficulty levels for the specific image editing task. The quality metrics—PSNR, SSIM, LPIPS, and FID—are calculated for each pair of watermarked and input images and then averaged across all 10,000 image pairs. MBRS and StegaStamp perform well in image regeneration and local editing tasks; however, they have lower image quality. Qualitative comparison is provided in Appendix H. Additionally, MBRS has a limited encoding capacity of only 30 bits. Although SepMark, PIMoG, and TrustMark strike a better balance between image quality and detection accuracy, their detection accuracy remains unsatisfactory. In contrast, our methods, VINE-B and VINE-R, offer the best trade-off. VINE-B provides superior image quality with slightly lower detection accuracy under image editing, while VINE-R delivers greater robustness by sacrificing some image quality.
@@
Figure 5 illustrates the watermarking performance across various difficulty levels for different image editing methods. The evaluation results against representative editing models are displayed, while additional results are available in Appendix I. For image regeneration, both VINE-R and MBRS maintain high TPR@0.1%FPR across various difficulty levels. VINE-B, StegaStamp, and PIMoG perform well when subjected to minor perturbations. It is important to note that stochastic regeneration with large noise perturbation steps can significantly alter the image, as shown in Appendix F.1. Although SSL achieves higher bit accuracy and AUROC than TrustMark, it exhibits a lower TPR@0.1%FPR. Further investigation revealed that SSL has a high FPR, often producing bit accuracy exceeding 0.7 for unwatermarked images. Therefore, bit accuracy and AUROC alone are insufficient for evaluating watermarking performance.
For global editing and local editing, VINE-B, VINE-R, and StegaStamp achieve high TPR@0.1%FPR across various classifier-free guidance scales. Notably, we also use UltraEdit for global editing, which provides better alignment between the edited image and the editing instructions compared to instruct-Pix2Pix and MagicBrush (a quantitative analysis of editing models is presented in Appendix F.2). However, methods that perform well for local editing with UltraEdit do not perform satisfactorily when applied to global editing using the same model, as shown in Figure 16(b). This suggests that global editing more significantly degrades watermarks. Image-to-video generation is not a form of traditional image editing, but we are interested in whether watermarks can persist in the generated frames. As illustrated in Figure 16(e), the overall detection rate is not high. Upon analyzing I2V generation in the frequency domain, we discovered that this process significantly reduces the intensity of nearly all patterns across all frequency bands, rendering them unobservable to the human eye. We infer this is because the generated video frames alter the image layout as objects or the camera move. In this case, the intensity of the watermarking patterns should be substantially increased, at least to levels exceeding those shown in Figure 6.

### v2

@@ Abstract
Current image watermarking methods are vulnerable to advanced image editing techniques enabled by large-scale text-to-image models. These models can distort embedded watermarks during editing, posing significant challenges to copyright protection. In this work, we introduce W-Bench, the first comprehensive benchmark designed to evaluate the robustness of watermarking methods against a wide range of image editing techniques, including image regeneration, global editing, local editing, and image-to-video generation. Through extensive evaluations of eleven representative watermarking methods against prevalent editing techniques, we demonstrate that most methods fail to detect watermarks after such edits. To address this limitation, we propose VINE, a watermarking method that significantly enhances robustness against various image editing techniques while maintaining high image quality. Our approach involves two key innovations: (1) we analyze the frequency characteristics of image editing and identify that blurring distortions exhibit similar frequency properties, which allows us to use them as surrogate attacks during training to bolster watermark robustness; (2) we leverage a large-scale pretrained diffusion model SDXL-Turbo, adapting it for the watermarking task to achieve more imperceptible and robust watermark embedding. Experimental results show that our method achieves outstanding watermarking performance under various image editing techniques, outperforming existing methods in both image quality and robustness. Code is available at https://github.com/Shilin-LU/VINE.
@@ 2 Related Work
Robust watermarking. Image watermarking has long been studied for purposes such as tracking and protecting intellectual property (Al-Haj, 2007; Cox et al., 2007; Navas et al., 2008). Recently, deep learning-based methods (Bui et al., 2023; Chen & Li, 2024; Fang et al., 2022; 2023; Jia et al., 2021; Kishore et al., 2021; Luo et al., 2020; 2024; Ma et al., 2022; Tancik et al., 2020; Wu et al., 2023; Zhu, 2018; Zhang et al., 2019; 2021), have demonstrated competitive robustness against a wide range of transformations. However, these methods remain vulnerable to image editing powered by large-scale generative models. Three recent studies—EditGuard (Zhang et al., 2024d), Robust-Wide (Hu et al., 2024), and JigMark (Pan et al., 2024)—have begun to develop watermarking models that are robust to such image editing. However, although EditGuard, which replaces the unedited regions of an edited image with corresponding areas of the unedited watermarked version, demonstrates good performance, its application in real-world scenarios could be ineffective if the detector has no access to the unedited watermarked image. In our benchmark, we assume that the detector has access only to the edited image without any additional information. JigMark employs contrastive learning to train a watermark encoder and a classifier that determines if an image is watermarked, but it does not support decoding free-form messages. Robust-Wide incorporates Instruct-Pix2Pix into the noise layer using gradient truncation, but its generalization to other editing models could be less effective. Moreover, neither JigMark nor Robust-Wide has been open-sourced.
@@ 3.1 Frequency nature of image editing
To develop a robust watermarking model against image editing, it should learn to embed information into the low-frequency bands. To identify effective surrogate attacks, we explore various image distortions, denoted as $\mathcal{T}(\cdot)$, that resemble image editing to some extent. Both image editing and distortion methods preserve the overall image layout and most of the content, although image distortions typically result in lower perceptual quality. Notably, as shown in Figure 3, certain blurring distortions (e.g., pixelation and defocus blur) exhibit a similar trend to image editing. In contrast, widely used distortions, such as JPEG compression and saturation, do not display this behavior. Since these blurring distortions are computationally efficient, we incorporate them with varying severity levels into our noise layers during training. This encourages the model to embed information within the low-frequency bands (see the frequency patterns of each watermarking method in Appendix B). Consequently, the robustness against image editing is enhanced, as demonstrated by the ablation study in Table 2. The complete set of distortions applied in the noise layer includes common distortions such as saturation, contrast, brightness adjustments, JPEG compression, Gaussian noise, shot noise, impulse noise, and speckle noise to counteract image degradation caused by transmission. Additionally, we incorporate various blurring distortions—such as pixelation, defocus blur, zoom blur, Gaussian blur, and motion blur—to resist image editing.
@@ 3.3 Objective function and training strategy
Training strategy. In the first training stage, we prioritize the watermark extraction loss by setting $\alpha$ to 10 and $\beta_{\text{MSE}}$, $\beta_{\text{LPIPS}}$, $\beta_{\text{GAN}}$ each to 0.01. To preserve the generative prior, the UNet and VAE decoder of the SDXL-Turbo, along with the added zero-convolution layers, are frozen. Once the bit accuracy exceeds 0.85, we transition to the second stage, unfreezing all parameters for further training. At this point, the loss weighting factors are adjusted to $\alpha=1.5$, $\beta_{\text{MSE}}=2.0$, $\beta_{\text{LPIPS}}=1.5$, and $\beta_{\text{GAN}}=0.5$. The model after the first two stages serves as our base model, dubbed VINE-B. We then fine-tune VINE-B in the third stage by incorporating Instruct-Pix2Pix (Brooks et al., 2023), a representative instruction-driven image editing model, into our noise layer. Gradients are backpropagated through a straight-through estimator (Bengio et al., 2013). Note that it cannot converge if it is applied directly during the early training stages. The fine-tuned model is referred to as VINE-R. Additional implementation details are provided in Appendix F.
Resolution scaling. Different watermarking models are typically trained using a fixed input resolution, limiting them to accepting only fixed-resolution inputs during testing. However, in practical applications, supporting watermarking at the original resolution is crucial to preserve the input image quality. Bui et al. (2023) propose a method (detailed in Appendix D.1) to adapt any watermarking model to handle arbitrary resolutions without compromising the quality of the watermarked images and their inherent robustness, as demonstrated in Appendices D.2 and D.3. In our experiments, we apply this resolution scaling method to all methods, enabling them to operate at a uniform resolution of $512\times 512$, which is compatible with image editing models.
@@ 4.1 Image editing methods
Image regeneration. Image regeneration involves perturbing an image into a noisy version and then reconstructing it. The perturbing process can be either stochastic or deterministic. In the stochastic method (Meng et al., 2021; Nie et al., 2022; Zhao et al., 2023b), random Gaussian noise is introduced to the image, with the noise level typically controlled by a timestep $t_{s}$ using common noise schedulers such as VP (Ho et al., 2020), VE (Song & Ermon, 2019; Song et al., 2020b), FM (Lipman et al., 2022; Liu et al., 2022a), and EDM (Karras et al., 2022). The diffusion model then denoises the noisy image starting from timestep $t_{s}$ to produce a clean image. In contrast, the deterministic method (Mokady et al., 2022; Song et al., 2020a; Wallace et al., 2022), also known as image inversion, utilizes a diffusion model to deterministicly invert a clean image into a noisy version through multiple sampling steps $n_{d}$. The image is then reconstructed by applying the same sampling methods over the same number of sampling steps $n_{d}$. We employ the widely used VP scheduler for stochastic regeneration, testing noise timesteps $t_{s}$ from 60 to 240 in increments of 20. For deterministic regeneration, we utilize the fast sampler DPM-solver (Lu et al., 2022a; b) and evaluate sampling steps $n_{d}$ of 15, 25, 35, and 45. See Appendix G.2 for the effects of image regeneration on images.
@@ 4.2 Experimental setup
Table 1: Comparison of watermarking performance in terms of watermarked image quality and detection accuracy across various image editing methods at a uniform resolution $512\times 512$. Quality metrics are averaged over 10,000 images, and the TPR@0.1%FPR for each specific editing method is averaged over 5,000 images. The best value in each column is highlighted in bold, and the second best value is underlined. Abbreviations: Cap = Encoding Capacity; Sto = Stochastic Regeneration; Det = Deterministic Regeneration; Pix2Pix = Instruct-Pix2Pix; Ultra = UltraEdit; Magic = MagicBrush; CtrlN = ControlNet-Inpainting; SVD = Stable Video Diffusion.
@@
Datasets. We train VINE using the OpenImage dataset (Kuznetsova et al., 2020) at a resolution of $256\times 256$. The training details are provided in Appendix F. For evaluation, we randomly sample 10,000 instances from the UltraEdit dataset (Zhao et al., 2024c), each containing a source image, an editing prompt, and a region mask. The images in UltraEdit dataset are photographs sourced from datasets such as COCO (Lin et al., 2014), Flickr (Young et al., 2014), and ShareGPT4V (Chen et al., 2023). Of these 10,000 samples, 1,000 are allocated for stochastic regeneration, another 1,000 for deterministic regeneration, and 1,000 for global editing. For local editing, 5,000 samples are designated for local editing. These consist of five sets, each containing 1,000 images, which are edited using mask sizes covering 10–20%, 20–30%, 30–40%, 40–50%, and 50–60% of the total image area. Additionally, we include 1,000 samples for image-to-video generation and 1,000 for testing conventional distortion, thereby completing the total evaluation set.
Baselines. We compare VINE with eleven watermark baselines, all utilizing their officially released checkpoints. These baselines include MBRS (Jia et al., 2021), CIN (Ma et al., 2022), PIMoG (Fang et al., 2022), RivaGAN (Zhang et al., 2019), SepMark (Wu et al., 2023), TrustMark (Bui et al., 2023), DWTDCT (Al-Haj, 2007), DWTDCTSVD (Navas et al., 2008), SSL (Fernandez et al., 2022), StegaStamp (Tancik et al., 2020), and EditGuard (Zhang et al., 2024d). Although the baselines were trained at different fixed resolutions (as detailed in Appendix D), we apply resolution scaling (described in Section 3.3) to standardize all of them to a uniform resolution of $512\times 512$. This standardization does not compromise their robustness, as demonstrated in Appendix D.3.
Metrics. We evaluate the imperceptibility of watermarking models using standard metrics, including PSNR, SSIM, LPIPS (Zhang et al., 2018), and FID (Parmar et al., 2022). For watermark extraction, it is essential to strictly control the false positive rate (FPR), as incorrectly labeling non-watermarked images as watermarked can be detrimental—a concern often overlooked in previous studies. Neither high bit accuracy nor AUROC alone guarantees a high true positive rate (TPR) at a low FPR. A detailed discussion is provided in Appendix C. Thus, we primarily focus on TPR@0.1%FPR and TPR@1%FPR as our main metrics. Accordingly, both the watermarked and original images are fed into watermark decoders for evaluation. Additionally, we also provide bit accuracy and AUROC for reference. Note that all reported baseline bit accuracies do not include error correction methods, such as BCH (Bose & Ray-Chaudhuri, 1960), which can be applied to all watermarking models.
@@ 4.3 Benchmarking results and analysis
Table 1 summarizes the overall evaluation results. As discussed in Section 4.2, we report TPR@0.1%FPR as the primary metric, with additional metrics provided in Figure 5. Each reported TPR@0.1%FPR value is averaged across a total of $m\times$1,000 images, where $m$ represents the number of difficulty levels for the specific image editing task. The quality metrics—PSNR, SSIM, LPIPS, and FID—are calculated for each pair of watermarked and input images and then averaged across all 10,000 image pairs. MBRS and StegaStamp perform well in image regeneration and local editing tasks; however, they have lower image quality. Qualitative comparison is provided in Appendix I. Additionally, MBRS has a limited encoding capacity of only 30 bits. Although SepMark, PIMoG, and TrustMark strike a better balance between image quality and detection accuracy, their detection accuracy remains unsatisfactory. In contrast, our methods, VINE-B and VINE-R, offer the best trade-off. VINE-B provides superior image quality with slightly lower detection accuracy under image editing, while VINE-R delivers greater robustness by sacrificing some image quality. EditGuard is not designed for robust watermarking against image editing, as it is trained with mild degradation. Instead, it offers a feature for tamper localization, enabling the identification of edited regions.
@@
Figure 5 illustrates the watermarking performance across various difficulty levels for different image editing methods. The evaluation results against representative editing models are displayed, while additional results are available in Appendix J. For image regeneration, both VINE-R and MBRS maintain high TPR@0.1%FPR across various difficulty levels. VINE-B, StegaStamp, and PIMoG perform well when subjected to minor perturbations. It is important to note that stochastic regeneration with large noise perturbation steps can significantly alter the image, as shown in Appendix G.2. Although SSL achieves higher bit accuracy and AUROC than TrustMark, it exhibits a lower TPR@0.1%FPR. Further investigation revealed that SSL has a high FPR, often producing bit accuracy exceeding 0.7 for unwatermarked images. Therefore, bit accuracy and AUROC alone are insufficient for evaluating watermarking performance.
For global editing and local editing, VINE-B, VINE-R, and StegaStamp achieve high TPR@0.1%FPR across various classifier-free guidance scales. Notably, we also use UltraEdit for global editing, which provides better alignment between the edited image and the editing instructions compared to instruct-Pix2Pix and MagicBrush (a quantitative analysis of editing models is presented in Appendix G.3). However, methods that perform well for local editing with UltraEdit do not perform satisfactorily when applied to global editing using the same model, as shown in Figure 18(b). This suggests that global editing more significantly degrades watermarks. Image-to-video generation is not a form of traditional image editing, but we are interested in whether watermarks can persist in the generated frames. As illustrated in Figure 18(e), the overall detection rate is not high. Upon analyzing I2V generation in the frequency domain, we discovered that this process significantly reduces the intensity of nearly all patterns across all frequency bands, rendering them unobservable to the human eye. We infer this is because the generated video frames alter the image layout as objects or the camera move. In this case, the intensity of the watermarking patterns should be substantially increased, at least to levels exceeding those shown in Figure 7.